How much pure hydrocloric acid must be added to 40 milliliters of a 50% solution to obtain a 75% acid solution?
Answer by stanbon(75887)

How much pure hydrocloric acid must be added to 40 milliliters of a 50% solution to obtain a 75% acid solution?
-----
Equation:
acid + acid = acid
0.50*40 + x = 0.75(40+x)
-------
Multiply thru by 100 to get:
50*40 + 100x = 75*40 + 75x
25x = 25*40
x = 40 milliters (amt. of pure HCL needed for the mixture)
===============================
Cheers,
Stan H.
===========